OpenStudy (anonymous):

Bought two plots of land for $120,000 made a profit of 15% on on plot. Lost 10% on the other plot. Total profit was $5500 how much did she pay for each property

OpenStudy (anonymous):

put x = amount paid for the property than gained 15% and so 120,000 - x is amount paid for the property that lost 10% then solve \[.15x-.10(120,000-x)=55,000\] for x

OpenStudy (anonymous):

got it from here? you might want to multiply both sides by 10 to get \[15x-10(120,000-x)=5,500,000\] \[25x-1,200,000=5,500,000\] \[25x=6,700,000\] and finally \[x=6,700,000\div 25\]